Let S = {(-6, 0, 3),(0, -7, -7),(0,2,0)} be an ordered basis of R3.
Let v be a vector in R3, v=(4,7,-1)

You calculate V in the basis of S. And get:

(a1, a1, a3)

What is the value of a3?
Answer #1

Date V= a. A t de B + Q3 C T-6 7 7 ar tas 2 o 3 we 4 = - 691 7 = fazt 203 1 3 al 792 Bolving lving above egun au= -2/3 az = -